OpenStudy (anonymous):

On Monday, a scientist starts with 3 cells in a dish. Each day, the number of cells triples. Which expression could be used to describe the number of cells in the dish on Wednesday? a) 3 b) 3^2 c) 3^3 d) 3^4

OpenStudy (anonymous):

(d) 3^4

OpenStudy (anonymous):

3^3 since you have 3 days which the cells triple.

OpenStudy (kirbykirby):

Monday: 3 cells Tuesday: 9 cells (3*3) Wednesday: 27 cells (9*3) = 3*3*3 = 3^3